(10000, -6000, -5500)

(10000, -3200, -8960)

(10000, -6800, -4320)

a. What is the interest rate for each of the cash flows?
Answer #1

a)

i)

Let r be interest rate

(10000, -6000, -5500)

NPV=10000-6000/(1+r)-5500/(1+r)^2

Put NPV=0

0=10000-6000/(1+r)-5500/(1+r)^2

10000*(1+r)^2-6000*(1+r)-5500=0

Let 1+r=x

10000x^2-6000x-5500=0

100x^2-60x-55=0

60 -b V04 - 4ac 2a (-60)2 - 4* 100 * (-55) 2 * 100

T = 60+ 160 = 1.1 and - 0.50 2 * 100

We know that 1+r=x

for x=1.1, r=0.1 or 10%

for x=-0.50, r=-1.50 or -150%

We ignore negative value of r,

So, interest rate=10%

ii)

Let r be interest rate

(10000, -3200, -8960)

NPV=10000-3200/(1+r)-8960/(1+r)^2

Put NPV=0

0=10000-3200/(1+r)-8960/(1+r)^2

10000*(1+r)^2-3200*(1+r)-8960=0

Let 1+r=x

10000x^2-3200x-8960=0

1000x^2-320x-896=0

-b + 320 + b2 - 4ac 2a (-320)2 - 4* 1000* (-896) 2 * 1000

= 320 = 1920 2 * 1000 = 1.12 and – 0.80

We know that 1+r=x

for x=1.12, r=0.12 or 12%

for x=-0.80, r=-1.80 or -180%

We ignore negative value of r,

So, interest rate=12%

iii)

Let r be interest rate

(10000, -6800, -4320)

NPV=10000-6800/(1+r)-4320/(1+r)^2

Put NPV=0

0=10000-6800/(1+r)-4320/(1+r)^2

10000*(1+r)^2-6800*(1+r)-4320=0

Let 1+r=x

10000x^2-6800x-4320=0

1000x^2-680x-432=0

- 6 680 62 - 4ac 2a -680-)2 - 4 * 1000 * (-432) 2 * 1000

T = 680 + 1480 2 * 1000 = 1.08 and - 0.40

We know that 1+r=x

for x=1.08, r=0.08 or 8%

for x=-0.40, r=-1.40 or -140%

We ignore negative value of r,

So, interest rate=8%